How much heat (in kJ) is required to warm 13.0 g of ice, initially at -14.0 ∘C, to steam at 110.0 ∘C? The heat capacity of ice is 2.09 J/g⋅∘C and that of steam is 2.01 J/g⋅∘C.
Answer #1

We require 2 type of heat, latent heat and sensible heat

Sensible heat (CP): heat change due to Temperature difference

Latent heat (LH): Heat involved in changing phases (no change of T)

Then

Q1 = m*Cp ice * (Tf – T1)

Q2 = m*LH ice

Q3 = m*Cp wáter * (Tb – Tf)

Q4 = m*LH vap

Q5 = m*Cp vap* (T2 – Tb)

Note that Tf = 0°C; Tb = 100°C, LH ice = 334 kJ/kg; LH water = 2264.76 kJ/kg.

Cp ice = 2.01 J/g°C ; Cp water = 4.184 J/g°C; Cp vapor = 2.030 kJ/kg°C

Then

Q1 = 13*2.09 * (0 – -14) = 380.38

Q2 = 13*334 = 4342

Q3 = 13*4.184 * (100 – 0) = 5439.2

Q4 = 13*2264.76 = 29441.88

Q5 = 13*2.01* (110– 100) = 261.3

QT = Q1+Q2+Q3+Q4+Q5 = 380.38 +4342+5439.2 + 29441.88+261.3 = 39864.76 J = 39.86 kJ